# Turán number of complete multipartite graphs in multipartite graphs

### v1

@@ Abstract.
In this paper we study a multi-partite version of the Erdős–Stone theorem. Given integers $r<k$ and $t\geq 1$, let $\operatorname{ex}_{k}(n,K_{r+1}(t))$ be the maximum number of edges of $K_{r+1}(t)$-free $k$-partite graphs with $n$ vertices in each part, where $K_{r+1}(t)$ is the $t$-blowup of $K_{r+1}$. An easy consequence of the supersaturaion result gives that $\operatorname{ex}_{k}(n,K_{r+1}(t))=\operatorname{ex}_{k}(n,K_{r+1})+o(n^{2})$. Similar to a result of Erdős and Simonovits for the non-partite case, we find that the error term is closely related to the (multi-partite) Zarankiewicz problem. Using such Zarankiewicz numbers, for $t=2,3$, we determine the error term up to an additive linear term; using some natural assumptions on such Zarankiewicz numbers, we determine the error term up to an additive constant depending on $k$, $r$ and $t$. We actually obtain exact results in many cases, for example, when $k\equiv 0,1\pmod{r}$. Our proof uses the stability method and starts by proving a stability result for $K_{r+1}$-free multi-partite graphs.
@@ 1. Introduction
Extremal problems whose host graphs are multipartite graphs have been studied since 1951, when Zarankiewicz proposed the study of the largest number of edges in a bipartite graph not containing a copy of $K_{s,t}$. Let $\mathcal{G}(n_{1},\dots,n_{k})$ denote the family of $k$-partite graphs with $n_{1},\dots,n_{k}$ vertices in its parts and write $\mathcal{G}_{k}(n)=\mathcal{G}(n,\dots,n)$ with $k$ parts. Given a graph $F$, define $\operatorname{ex}(n_{1},\dots,n_{k};F)$ as the largest number of edges in $F$-free graphs from $\mathcal{G}(n_{1},\dots,n_{k})$, and let $\operatorname{ex}_{k}(n,F)=\operatorname{ex}(n,\dots,n;F)$ (with $k$ parts). (Trivially $\operatorname{ex}_{k}(n,F)=\binom{k}{2}n^{2}$ if the chromatic number $\chi(F)>k$.) In 1975 Bollobás, Erdős, and Szemerédi [MR0389639] investigated several Turán-type problem for multipartite graphs. Applying a simple counting argument, they showed that
@@
for any $n,k,r\in\mathbb{N}$ with $k\geq r$. The main results of [MR0389639] concern the minimum degree version of this problem, which has been intensively studied, see [MR2195582, MR1860440, MR1208805, MR4496024, MR2246152].
In this paper we study $\operatorname{ex}_{k}(n,K_{r+1}(t))$, the multi-partite versions of the Erdős–Stone theorem and Theorem 1.
@@
We first prove a stability theorem for $\operatorname{ex}_{k}(n,K_{r+1})$. This result was independently obtained by Chen, Lu, and Yuan [CLY23]. Given $r,k\in\mathbb{N}$ with $k\geq r$, write $k=ar+b$ for $10\leq b\leq r-$ 1. By Turán’s theorem, the Turán graph $T_{r}(k)=K_{a,\dots,a,a+1,\dots,a+1}$ (with $b$ parts of size $a+1$ and $r-b$ parts of size $a$) is the unique largest $K_{r+1}$-free graph on $k$ vertices. We now describe extremal graphs for $\operatorname{ex}_{k}(n,K_{r+1})$. Let ${\mathcal{T}}_{r,k}(n)$ be the collection of $k$-partite graphs with parts $V_{1},\dots,V_{k}$ of size $n$ defined as follows. If $b>0$, we arbitrarily divide $V_{ar+1},\dots,V_{k}$ into $r$ sets $W_{1},\dots,W_{r}$ (some of them may be empty) such that each $W_{i}$ is a subset of $V_{j}$ for some $j$; if $b=0$, then let $W_{1},\dots,W_{r}$ be empty sets. Now let $T$ be the $r$-partite graph with parts $U_{1},\dots,U_{r}$ such that
@@
obtained from the complete $r$-partite graph $K(U_{1},\dots,U_{r})$ by removing edges between $W_{i}$ and $W_{i^{\prime}}$, $i\neq i^{\prime}$, whenever $W_{i},W_{i^{\prime}}\subseteq V_{j}$ for some $j$ (in other words, $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$). Since $T$ is $r$-partite, it is $K_{r+1}$-free. Let $U=\bigcup_{i\in[r]}U_{i}$ and $W=\bigcup_{i\in[r]}W_{i}$. Note that $e_{T}(U)=\binom{r}{2}a^{2}n^{2}$ and $e_{T}(W,T)=|W|(k-a-1)n=b(k-a-1)n^{2}$. Since $t_{r}(k)=\binom{r}{2}a^{2}+b(k-a-1)$, it follows that $e(T)=t_{r}(k)n^{2}$. By (1.1), $T$ is an extremal graph for $K_{r+1}$. Let ${\mathcal{T}}_{r,k}(n)$ be the collection of all such $T$.
Given two graphs $G,H\in\mathcal{G}_{k}(n)$ on the same parts $V_{1},\dots,V_{k}$, we say that $G$ and $H$ are $\gamma$-close if $|E(G)\triangle E(H)|\leq\gamma n^{2}$.
Theorem 2.
For any positive integers $r\leq k$ and any $\gamma>0$, there exist $\varepsilon>0$ and $n_{0}$ such that the following holds for every integer $n\geq n_{0}$. Suppose $G\in\mathcal{G}_{k}(n)$ is $K_{r+1}$-free and $e(G)\geq(t_{r}(k)-\varepsilon)n^{2}$. Then $G$ is $\gamma$-close to a member of ${\mathcal{T}}_{r,k}(n)$.
We now consider $\operatorname{ex}_{k}(n,K_{r+1}(t))$, We assume that $k\geq r+1$ because otherwise $\text{ex}_{k}(n,K_{r+1}(t))=\binom{k}{2}n^{2}$ trivially. Applying (1.1) and either the Regularity Lemma or the Graph Removal Lemma, one can easily derive the following Erdős–Stone theorem for multipartite graphs:
$\text{ex}_{k}(n,K_{r+1}(t))=t_{r}(k)n^{2}+o(n^{2}).$
Applying Theorem 2 and the Graph Removal Lemma, we derive the following stability result for $\operatorname{ex}_{k}(n,K_{r+1}(t))$, which handles the non-extremal case for $\operatorname{ex}_{k}(n,K_{r+1}(t))$.

### v2

@@ Abstract.
In this paper we study a multi-partite version of the Erdős–Stone theorem. Given integers $r<k$ and $t\geq 1$, let $\operatorname{ex}_{k}(n,K_{r+1}(t))$ be the maximum number of edges of $K_{r+1}(t)$-free $k$-partite graphs with $n$ vertices in each part, where $K_{r+1}(t)$ is the complete $(r+1)$-partite graph with $t$ vertices in each part. We determine the exact value of $\operatorname{ex}_{k}(n,K_{r+1}(t))$ for $t\leq 3$, $r<k\leq 2r$ and sufficiently large $n$. We also characterize all extremal graphs for $r,k$ such that $r$ divides $k$, analogous to a result of Erdős and Simonovits on forbidding $K_{r+1}(t)$ in general graphs.
@@ 1. Introduction
Extremal problems with multipartite graphs as host graphs have been studied since 1951, when Zarankiewicz proposed the study of the largest number of edges in a $K_{s,t}$-free bipartite graph. Let $\mathcal{G}(n_{1},\dots,n_{k})$ denote the family of $k$-partite graphs with $n_{1},\dots,n_{k}$ vertices in its parts and write $\mathcal{G}_{k}(n)=\mathcal{G}(n,\dots,n)$ with $k$ parts. Given a graph $F$, define $\operatorname{ex}(n_{1},\dots,n_{k};F)$ as the largest number of edges in $F$-free graphs from $\mathcal{G}(n_{1},\dots,n_{k})$, and let $\operatorname{ex}_{k}(n,F)=\operatorname{ex}(n,\dots,n;F)$ (with $k$ parts). (Trivially $\operatorname{ex}_{k}(n,F)=\binom{k}{2}n^{2}$ if the chromatic number $\chi(F)>k$.) In 1975 Bollobás, Erdős, and Szemerédi [MR0389639] investigated several Turán-type problem for multipartite graphs. Applying a simple counting argument, they showed that, for any $n,k,r\in\mathbb{N}$ with $k>r$,
@@
The main results of [MR0389639] were on the minimum degree that forces a copy of $K_{r}$ in the graphs of $\mathcal{G}_{k}(n)$. This problem has been intensively studied and resolved when $k=r$ (frequently in its complementary form concerning independent transversals) [MR2195582, MR1860440, MR1208805, MR2246152], and more recently for $k>r$ [MR4496024] (certain cases are still unsolved). There are several other extremal results for multipartite graphs. Bollobás, Erdős, and Straus [MR0379256] determined $\operatorname{ex}(n_{1},\dots,n_{r};K_{r})$ for all $n_{1},\dots,n_{r}$ and $r$. Let $kK_{r}$ denote $k$ vertex-disjoint copies of $K_{r}$. The problem $\operatorname{ex}(n_{1},\dots,n_{r};kK_{r})$ were studied more recently [MR3724116, MR4425157, MR4595892] and settled by Chen, Lu, and Yuan [CLY23] when $n_{1},\dots,n_{r}$ are sufficiently large ($k,r$ are arbitrary but fixed). The minimum pair density of multipartite graphs that forces a clique was studied by Bondy, Shen, Thomassé, and Thomassen [bondy] and Pfender [pfender].
In this paper we study $\operatorname{ex}_{k}(n,K_{r+1}(t))$, the multi-partite version of the Erdős–Stone theorem and Theorem 1. To give the precise value of $\operatorname{ex}_{k}(n,K_{r+1}(t))$, we need the following definition. Given $a,t,n_{1},\dots,n_{a}\in\mathbb{N}$, let $z_{t}^{(a)}(n_{1},\dots,n_{a})$ be the $a$-partite Zarankiewicz number for $K_{t,t}$, that is, the maximum number of edges in a $K_{t,t}$-free bipartite graph with part sizes $n_{1},\dots,n_{a}$. For simplicity, we write $z_{t}^{(a)}(n)$ if $n_{1}=\cdots=n_{a}$. We also write $z_{t}(m,n)$ for $z_{t}^{(2)}(m,n)$.
Theorem 1 says that any extremal graph for $K_{r+1}(t)$ is the join of an extremal graph for $K_{t,t}$ and $r-1$ extremal graphs for $K_{1,t}$. Inspired by this, a natural guess of an extremal graph for its $r$-partite analogue is as follows. Suppose $k=ar+b$ with $0\leq b<r$. We start with an $n$-blowup of $T_{r}(k)$, which has $r$ classes and each class has either $a$ or $a+1$ parts. We add a $K_{t,t}$-free graph to one class with the most number of parts and $\{K_{1,t},K_{2,2}\}$-free graphs to the remaining $r-1$ classes. If $r<k\leq 2r$, then it is easy to see (as in the proof of Theorem 2) that this graph is $K_{r+1}(t)$-free and has $t_{r}(k)n^{2}+z_{t}(n,n)+(k-r-1)(t-1)n$ edges, and therefore
$\operatorname{ex}_{k}(n,K_{r+1}(t))\geq t_{r}(k)n^{2}+z_{t}(n,n)+(k-r-1)(t-1)n.$
In this paper we first improve this lower bound when $t\geq 2$ and $r<k\leq 2r$.
Theorem 2.
Suppose $t\geq 2$, $r<k\leq 2r$ and $n\geq 8t^{2}$. Then $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq g(n,r,k,t)$, where
$g(n,r,k,t):=t_{r}(k)n^{2}+z_{t}(n,n)+(t-1)(k-r-1)n+\min\{k-r-1,2r-k\}\left\lfloor\frac{(t-1)^{2}}{4}\right\rfloor.$
The following is our main result, in which we prove a matching upper bound when $t\in\{2,3\}$ and $n$ is sufficiently large.
Theorem 3.
For any $t\in\{2,3\}$ and integers $r$, there exists $n_{0}=n_{0}(r)\in\mathbb{N}$ such that for $n\geq n_{0}$ and $r<k\leq 2r$, we have $\operatorname{ex}_{k}(n,K_{r+1}(t))=g(n,r,k,t)$.
In fact, we conjecture that $\operatorname{ex}_{k}(n,K_{r+1}(t))=g(n,r,k,t)$ holds for all $t\geq 2$, $r<k\leq 2r$, and sufficiently large $n$.
A natural question is whether our result can be extended to larger values of $t$ and $k$. For larger value of $t$, although we can use $z_{t}(n,n)$ without knowing its precise value, we need several properties of this function in our proof. Kővári, Sós, Turán [MR0065617] showed that $z_{t}(n,n)=O(n^{2-1/t})$ for $t\geq 2$ and proving a matching lower bound is a well-known open problem:
  1. ($Z$)
    $z_{t}(n,n)=\Omega(n^{2-1/t})$ for $t\geq 2$.
It was shown [MR0200182, MR0223262] that ($Z$) holds for $t=2,3$. In addition, we will need the following properties.
  1. ($E1$)
    For any $a\in\mathbb{N}$, there exists $\delta>0$ such that for large $n$, $z_{t}^{(a+1)}(n)-z_{t}^{(a)}(n)\geq{\delta n^{2-1/t}}.$
  2. ($E2$)
    for any $\varepsilon\in(0,1]$, there exists $\delta>0$ such that for large $n$,
    $z_{t}(n,n)-z_{t}((1-\varepsilon)n,n)\geq{\delta n^{2-1/t}}.$
  3. ($E3$)
    $z_{t}(m,n)-z_{t}(m-1,n)\geq t-1$.
We can easily verify ($E1$)($E3$) for $t=2,3$. First, a proof of ($E1$) from ($Z$) for all $t$ is given in Section 3.3. Second, when $t=2,3$, ($E2$) follows from $z_{t}(m,n)\leq(1+o(1))mn^{1-1/t}$ by Füredi [MR1395691] and $z_{t}(n,n)\geq(1-o(1))n^{2-1/t}$ for $t=2,3$. Third, ($E3$) holds trivially because adding a vertex with $t-1$ edges to a $K_{t,t}$-free graph will not create a copy of $K_{t,t}$.
We do not know whether similar properties hold when $k>2r$: in ($E2$) we must deal with the $\lceil k/r\rceil$-partite Zarankiewicz number; we also need to replace $t-1$ by $\Omega(a^{2}t)$ in ($E3$), which seems out of reach at present.
Theorems 2 and 3 show that our problem is more complex than its non-partite counterpart, Theorem 1. Finally, we show that, when $r$ divides $k$, this additional complexity does not exist, and we give an analogue of Theorem 1, modulo the existence of a set of constantly many exceptional vertices.
Theorem 4.
For $r,k\in\mathbb{N}$ with $r\mid k$ and $t=2,3$, there exist $C_{0},n_{0}\in\mathbb{N}$ such that the following holds for $n\geq n_{0}$. Let $G$ be a $K_{r+1}(t)$-free $k$-partite graph with $n$ vertices in each part and $\operatorname{ex}_{k}(n,K_{r+1}(t))$ edges. Then there is a vertex partition of $G$ into $U_{1},\dots,U_{r}$, each consisting of exactly $k/r$ parts of $G$, and a vertex set $Z\subseteq V(G)$ with $|Z|\leq C_{0}$ such that
  • $G[U_{i}\setminus Z,U_{j}\setminus Z]$ is almost complete for all $i\neq j$,
  • $G[U_{1}\setminus Z]$ is $K_{t,t}$-free, and
  • $G[U_{2}\setminus Z],\dots,G[U_{r}\setminus Z]$ are $K_{1,t}$-free.
Showing $Z=\emptyset$ in Theorem 4 requires an $a$-partite analogue of ($E2$), which is unknown for $a\geq 3$.
For the rest of this paper we only consider $r\geq 2$, as it is easy to see that Theorems 24 hold for $r=1$.
Notation. Given integers $n\geq m\geq 1$, let $[n]=\{1,\dots,n\}$ and $[m,n]=\{m,m+1,\dots,n\}$. We omit floors and ceilings unless they are crucial, e.g., we may choose a set of $\varepsilon n$ vertices even if our assumption does not guarantee that $\varepsilon n$ is an integer.
2. Proof of Theorem 2
In this section we prove Theorem 2, that is, $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq g(n,r,k,t)$ for $r<k\leq 2r$. Our proof needs a $t$-regular $K_{2,2}$-free bipartite graph with $n$ vertices in each part. It is well known (see [MR101250]) that such graph exists for infinitely many $n\in\mathbb{N}$ with $n\geq t^{2}$. The following proposition from [MR2519934]*Section 2 allows $n$ to be any integer that is at least $8t^{2}$.
Proposition 2.1.
[MR2519934] For $t\geq 1$ and $n\geq 8t^{2}$, there exists a $t$-regular $K_{2,2}$-free bipartite graph with $n$ vertices in each part.
Now we prove our lower bound on $\text{ex}_{k}(n,K_{r+1}(t))$ stated in Theorem 2.
Proof of Theorem 2.
First assume $r<k\leq 2r$ and $t\geq 2$. Let $V_{1},\dots,V_{k}$ be disjoint sets of size $n$ and, if $k<2r$, let $V_{k+1},\dots,V_{2r}$ be empty sets. Let $G^{\prime}$ be a complete $r$-partite graph with parts $V_{1}\cup V_{r+1},V_{2}\cup V_{r+2},\dots,V_{r}\cup V_{2r}$. Moreover, we add to $G^{\prime}$ a maximum $K_{t,t}$-free bipartite graph with bipartition $V_{1}\cup V_{r+1}$, and a $(t-1)$-regular $K_{2,2}$-free bipartite graph on $V_{i}\cup V_{i+r}$ for $2\leq i\leq k-r$ (the existence of such graph is guaranteed by Proposition 2.1). The resulting graph is $K_{r+1}(t)$-free because a copy of $K_{r+1}(t)$ has at most $12t-$ 1 vertices in $V_{1}\cup V_{r+1}$ and at most $t$ vertices in $V_{i}\cup V_{i+r}$ for $2\leq i\leq k-r$. This graph has $t_{r}(k)n^{2}+z_{t}(n)+(t-1)(k-r-1)n$ edges, and thus, $\operatorname{ex}_{k}(n,K_{r+1}(t))\geq t_{r}(k)n^{2}+z_{t}(n)+(t-1)(k-r-1)n$. This proves the theorem when $t=2$ or $k=2r$.
Now assume $r<k<2r$ and $t\geq 2$. Let $b=k-r$. Our goal is to give a better construction that shows
$\operatorname{ex}_{k}(n,K_{r+1}(t))\geq t_{r}(k)n^{2}+z_{t}(n,n)+(t-1)(b-1)n+\min\{b-1,r-b\}\left\lfloor\frac{(t-1)^{2}}{4}\right\rfloor.$
Let $V_{i,j}$, $(i,j)\in[r]\times[2]$ be vertex sets, where $V_{b+1,2},\dots,V_{r,2}$ are empty sets and other sets have size $n$. Let $G=K(V_{1,1}\cup V_{1,2},\dots,V_{r,1}\cup V_{r,2})$ be the complete $r$-partite graph with parts $V_{1,1}\cup V_{1,2},\dots,V_{r,1}\cup V_{r,2}$. Thus $e(G)=t_{r}(k)n^{2}$.
We first revise the partition as follows. Let $t^{\prime}:=\lceil(t-1)/2\rceil$ and $b^{\prime}:=\min\{b-1,r-b\}$. Let $\{V_{i,j}^{\prime},(i,j)\in[r]\times[2]\}$ be obtained from $\bigcup V_{i,j}$ by moving a set $S_{i,1}$ of $t^{\prime}$ vertices from $V_{i,1}$ to $V_{i+b-1,1}$, and moving a set $S_{i,2}$ of $t^{\prime}$ vertices from $V_{i,2}$ to $V_{i+b-1,2}$, for every $i\in[2,b^{\prime}+1]$. For $i\in[r]$, let $U_{i}:=V_{i,1}^{\prime}\cup V_{i,2}^{\prime}$ and $H:=K(U_{1},\dots,U_{r})\cap K(V_{1,1},\dots,V_{r,1},V_{1,2},\dots,V_{r,2})$. Let $H^{\prime}$ be obtained from $H$ by adding
  • a $K_{t,t}$-free bipartite graph on $U_{1}$ of size $z_{t}(n,n)$, and
  • a maximum $\{K_{1,t},K_{2,2}\}$-free bipartite graph on $U_{i}$ for $i\in[2,b]$.
  • $2t^{\prime}$ vertex-disjoint copies of $K_{1,t-1}$ on each $U_{i}$, $i\in[b+1,b+b^{\prime}]$, where the centers of stars are the $2t^{\prime}$ vertices of $S_{i-b+1,1}\cup S_{i-b+1,2}$; then add a copy of $K_{t^{\prime},t^{\prime}}$ on $S_{i-b+1,1}\cup S_{i-b+1,2}$. See Figure 1.
Figure 1. The lower bound construction for $K_{r+1}(t)$ with $r=5$, $k=8$ and $t=3$. The left figure is the standard construction similar to the one given in Theorem 1; the right figure is the construction presented in our proof of Theorem 3.
For $i\in[2,b]$, Proposition 2.1 implies that each $H^{\prime}[U_{i}]$ is $(t-1)$-regular and thus for $i\in[2,b^{\prime}+1]$, $H^{\prime}[U_{i}]$ has $(n-t^{\prime})(t-1)$ edges, and for $i\in[b^{\prime}+2,b]$ it has $n(t-1)$ edges; for $i\in[b+1,b+b^{\prime}]$, each $H^{\prime}[U_{i}]$ has $2t^{\prime}(t-1)+(t^{\prime})^{2}$ edges. Therefore, the number of edges in $H^{\prime}\setminus H$ is
$z_{t}(n,n)+(n-t^{\prime})(t-1)b^{\prime}+n(t-1)(b-b^{\prime}-1)+b^{\prime}(2t^{\prime}(t-1)+(t^{\prime})^{2})$
$=$ $z_{t}(n,n)+(t-1)(b-1)n+b^{\prime}(t^{\prime})^{2}+b^{\prime}t^{\prime}(t-1)$
$=$ $z_{t}(n,n)+(t-1)(b-1)n+2b^{\prime}(t^{\prime})^{2}+b^{\prime}t^{\prime}(t-1-t^{\prime})$
$=$ $z_{t}(n,n)+(t-1)(b-1)n+2b^{\prime}(t^{\prime})^{2}+\left\lfloor\frac{(t-1)^{2}}{4}\right\rfloor.$
We claim that $H$ contains $t_{r}(k)n^{2}-2b^{\prime}(t^{\prime})^{2}$ edges. Indeed, for every $i\in[2,b^{\prime}+1]$, the vertices of $S_{i,1}$ moved from $V_{i,1}$ to $V_{i+b-1,1}$ lose $n$ edges to $V_{i+b-1,1}$ and gain $n-t^{\prime}$ edges to $V_{i,2}$, thus having a net loss $(t^{\prime})^{2}$ edges between $U_{i}$ and $U_{i+b-1}$; the same holds for $S_{i,2}$. Thus, our claim holds after summing over all $b^{\prime}$ such rows, which implies that $e(H^{\prime})=t_{r}(k)n^{2}+z_{t}(n,n)+(t-1)(b-1)n+b^{\prime}\lfloor\frac{(t-1)^{2}}{4}\rfloor$.
At last, we show that $H^{\prime}$ is $K_{r+1}(t)$-free. Recall that by construction, every $U_{i}$ is triangle-free, $U_{1}$ is $K_{t,t}$-free, $U_{2},\dots,U_{b}$ are $\{K_{1,t},K_{2,2}\}$-free, and each of $U_{b+1},\dots,U_{U_{b+b^{\prime}}}$ induces vertex-disjoint copies of $K_{1,t-1}$ whose centers are joined by copies of $K_{t^{\prime},t^{\prime}}$. Suppose $K$ is a copy of $K_{r+1}(t)$ in $H^{\prime}$. By construction, $K$ contains at most $12t-$ 1 vertices in $U_{1}$. We claim that $|V(K)\cap(U_{i}\cup U_{i+b-1})|\leq 2t$ for $i\in[2,b^{\prime}+1]$ and $|V(K)\cap U_{j}|\leq t$ for $i\in[b^{\prime}+2,b]\cup[b+b^{\prime}+1,r]$, which will lead to a contradiction with $|V(K)|=(r+1)t$. Indeed, for $i\in[2,r]$, if $K$ contains at least $t+1$ vertices in $U_{i}$, then these vertices induce either a copy of $K_{1,t}$ or $K_{2,2}$. By construction, this is only possible for $i\in[b+1,b+b^{\prime}]$ and in that case $V(K)\cap U_{i}$ must intersect both $S_{i-b+1,1}$ and $S_{i-b+1,2}$. Furthermore, there exists $v\in V(K)\cap S_{i-b+1,1}$ and $v^{\prime}\in V(K)\cap S_{i-b+1,2}$ such that $v$ and $v^{\prime}$ are in different color classes of $K$. Since $v$ and $v^{\prime}$ have no common neighbor in $U_{i}\cup U_{i-b+1}$, $K$ induces at most two classes on $U_{i}\cup U_{i-b+1}$. Thus, $|V(K)\cap(U_{i}\cup U_{i-b+1})|\leq 2t$ and we are done. ∎
Remark. Note that for $t=2$, both our constructions give the same value, which means that the extremal graph is not unique. Indeed, it is easy to see that one can construct $b^{\prime}+1$ different ones – as we can move vertices for a subset of the $b^{\prime}$ rows.
3. Stability and proof outline
Let us first consider extremal graph for $\operatorname{ex}_{k}(n,K_{r+1})$. Given $r,k\in\mathbb{N}$ with $k>r$, write $k=ar+b$ for $10\leq b\leq r-$ 1. By Turán’s theorem, the Turán graph $T_{r}(k):=K_{a,\dots,a,a+1,\dots,a+1}$ (with $b$ parts of size $a+1$ and $r-b$ parts of size $a$) is the unique largest $K_{r+1}$-free graph on $k$ vertices. The following definition shows that there are many extremal graphs for $\operatorname{ex}_{k}(n,K_{r+1})$.
Definition 3.1.
@@ 3. Stability and proof outline
Let ${\mathcal{T}}_{r,k}(n)$ be the collection of $k$-partite graphs with parts $V_{1},\dots,V_{k}$ of size $n$ defined as follows. If $b>0$, we arbitrarily divide $V_{ar+1},\dots,V_{k}$ into $r$ sets $W_{1},\dots,W_{r}$ (some of them may be empty) such that each $W_{i}$ is a subset of $V_{j}$ for some $j$; if $b=0$, then let $W_{1},\dots,W_{r}$ be empty sets. Now let $T$ be the $r$-partite graph with parts $U_{1},\dots,U_{r}$ such that
@@
obtained from the complete $r$-partite graph $K(U_{1},\dots,U_{r})$ by removing edges between $W_{i}$ and $W_{i^{\prime}}$, $i\neq i^{\prime}$, whenever $W_{i},W_{i^{\prime}}\subseteq V_{j}$ for some $j$ (in other words, $T=K(U_{1},\dots,U_{r})\cap K(V_{1},\dots,V_{k})$).
Since $T$ is $r$-partite, it is $K_{r+1}$-free. Let $U=\bigcup_{i\in[r]}U_{i}$ and $W=\bigcup_{i\in[r]}W_{i}$. Note that $e_{T}(U)=\binom{r}{2}a^{2}n^{2}$ while the number of edges of $T$ incident to $W$ is equal to $|W|(k-a-1)n=b(k-a-1)n^{2}$. Since $t_{r}(k)=\binom{r}{2}a^{2}+b(k-a-1)$, it follows that $e(T)=t_{r}(k)n^{2}$. By (1.1), $T$ is an extremal graph for $K_{r+1}$. [^1] Indeed, Chen, Lu, and Yuan [CLY23] showed that ${\mathcal{T}}_{r,k}(n)$ are all the extremal graphs for $\operatorname{ex}_{k}(n,K_{r+1})$ and described all extremal graphs for $\operatorname{ex}(n_{1},\dots,n_{k};K_{r+1})$.
3.1. A stability theorem
We need the following stability result for $\operatorname{ex}_{k}(n,K_{r+1}(t))$. Given two graphs $G,H\in\mathcal{G}_{k}(n)$ on the same parts $V_{1},\dots,V_{k}$, we say that $G$ and $H$ are $\gamma$-close if $|E(G)\triangle E(H)|\leq\gamma n^{2}$.